# Dual-Unitary Classical Shadow Tomography

### v2

@@ Abstract
We introduce a classical shadow tomography scheme based on dual-unitary brick-wall circuits termed “dual-unitary shadow tomography” (DUST). For this we study operator spreading and Pauli weight dynamics in one-dimensional qubit systems, evolved by random two-local dual-unitary gates arranged in a brick-wall structure, ending with a final measurement layer. We do this by deriving general constraints on the Pauli weight transfer matrix and specializing to the case of dual-unitarity. We first show that dual-unitaries must have a minimal amount of entropy production. Remarkably, we find that operator spreading in these circuits have a rich structure resembling that of relativistic quantum field theories, with massless chiral excitations that can decay or fuse into each other, which we call left- or right-movers. We develop a mean-field description of the Pauli weight in terms of $\rho(x,t)$, which represents the probability of having nontrivial support at site $x$ and depth $t$ starting from a fixed weight distribution. We develop an equation of state for $\rho(x,t)$, and simulate it numerically using Monte Carlo simulations. Lastly, we demonstrate that the fast-thermalizing properties of dual-unitary circuits make them better at predicting large operators than shallow brick-wall Clifford circuits. Our results are robust to finite-size effects due to the chirality of dual-unitary brick-wall circuits.
@@ I Introduction
We begin by giving a brief overview of classical shadow tomography as it is relevant to this paper. Next, we introduce the concept of dual-unitarity in circuits and our notion of random, locally-scrambled dual-unitary gates. Equipped with this knowledge, we motivate the setup and the Pauli weight transfer matrix which defines how the Pauli weight evolves as a state vector. Using the definitions of dual-unitary and some information-theoretic bounds, we derive the general form of the transfer matrix in the dual-unitary case. Then, we study the dynamics implied by this transfer matrix as a set of Feynman rules on the left and right-movers. Essentially, the movers are allowed to fuse and by emitted by one another, but one cannot arbitrarily transform into the other. This is the key difference with transfer matrices which are derived form two-designs. We then introduce the average occupation $\rho(x,t)$, which characterizes the average Pauli weight on a single site, and we study its structure via Monte Carlo simulations. Then, we derive a mean-field equation for the dynamics of the average occupation, argue for why this description remains accurate, analyze the equation of state, and solve it in the homogeneous case. Lastly, we argue how these dual-unitary brick-wall circuits are advantageous compared to Clifford brick-wall circuits for the prediction of extensive observables using classical shadow tomography. This introduces our dual-unitary shadow tomography (DUST) scheme.
@@ II.1.1 General Formalism
is the locally-scrambled shadow norm $\|P\|_{\mathcal{E}_{\sigma}}^{2}$, which characterizes the sample complexity of predicting $P$ using the tomography scheme defined by $\mathcal{E}_{\sigma}$ [42, 48].
@@ II.1.2 Dual-Unitarity in Qubits
$V(J)=\begin{pmatrix}e^{iJ}&0&0&0\\ 0&0&-ie^{iJ}&0\\ 0&-ie^{iJ}&0&0\\ 0&0&0&e^{iJ}\end{pmatrix}.$ (7)
At $J=\pi/4$, $V$ is proportional to a swap gate and simply interchanges sites, scrambling no information. As we turn $J$ away from $\pi/4$, we scramble more and more information until we reach $J=0$ which is akin to Hamiltonian generated scrambling with quantum $XY$-model.
@@
In practice, it suffices to sample the local scramblers $u_{\pm},v_{\pm}$ according to the Clifford group ${\mathcal{C}\ell}(2)$. This is because ${\mathcal{C}\ell}(2)$ forms a 3-design, and since the tomographic properties of the measurement channel (e.g. the Pauli weight, the shadow norm, etc.) only depend on the second moment of the unitary distribution, all such quantities will be the same.
II.1.3 Set up
@@
II.2 Pauli Weight Transfer Matrix
@@
II.3 Dual-unitary case
@@
II.4 Pauli Weight Dynamics
@@ II.5 Occupation distribution
To refine our analysis, recall that the Pauli weight defines a probability distribution over all possible sub-regions of $\Omega=\{1\dots n\}$, or equivalently, all possible $n$-bit strings (see Eq. (18). This is because the transfer matrix which evolves the Pauli weight vector defined in Sec. II.1.3 is a probability-conserving process. Therefore, we have that the weights are positive i.e. $w_{P,t}(A):={\langle b|w_{P,t}\rangle}>0$, where $b$ is related to $A$ by Eq. (18), and also that
@@
II.6 Mean Field Approximation of Pauli Weight Dynamics
@@
II.7 Ramifications for the shadow norm and prediction
@@
which decays exponentially with a base of $e^{\log(25/16)}$, compared to $e^{\alpha^{\prime}}$ in the dual-unitary case. It follows that the dual-unitary should approach the Clifford limit more quickly provided that $\alpha^{\prime}>\log(25/16)$, which occurs around $\alpha\approx 0.35$. However, we caution that in practice, it may require a very deep circuit to see the advantage for $\alpha<2/3$.
@@ IV Acknowledgements
Appendix A Deriving the shadow norm bound via the mean-field approximation

### v3

# Dual-unitary shadow tomography

@@ Abstract
We introduce a classical shadow tomography protocol based on dual-unitary brick-wall circuits termed “dual-unitary shadow tomography” (DUST). To quantify the performance of DUST, we study operator spreading and Pauli weight dynamics in one-dimensional qubit systems, evolved by random two-local dual-unitary gates arranged in a brick-wall structure, ending with a final measurement layer. We do this by deriving general constraints on the Pauli weight transfer matrix and specializing to the case of dual-unitarity. Remarkably, we find that operator spreading in these circuits have a rich structure resembling that of relativistic quantum field theories, with massless chiral excitations that can decay or fuse into each other, which we call left- or right-movers. We develop a mean-field description of the Pauli weight in terms of $\rho(x,t)$, which represents the probability of having nontrivial support at site $x$ and depth $t$ starting from a fixed weight distribution. We develop an equation of state for $\rho(x,t)$ and simulate it numerically using Monte Carlo simulations. For the task of predicting operators with (nearly) full support, we show that DUST outperforms brick-wall Clifford shadows of equivalent depth. This advantage is further pronounced for small system sizes and more generally, our results are robust to finite-size effects.
@@ I Introduction
We begin by giving a brief overview of classical shadow tomography as it is relevant to this paper in Section II.1.1. Next, we introduce the concept of dual-unitarity in circuits and our notion of random, locally-scrambled dual-unitary gates in Section II.1.2. Equipped with this knowledge, we motivate the setup and the Pauli weight transfer matrix which defines how the Pauli weight evolves as a state vector in Section II.3. Using the definitions of dual-unitary and some information-theoretic bounds, we derive the general form of the transfer matrix in the dual-unitary case in Section II.4. Then, we study the dynamics implied by this transfer matrix as a set of Feynman rules on the left and right-movers. Essentially, the movers are allowed to fuse and by emitted by one another, but one cannot arbitrarily transform into the other. This is the key difference with transfer matrices which are derived form two-designs as discussed in Section II.5. We then introduce the average occupation $\rho(x,t)$, which characterizes the average Pauli weight on a single site, and we study its structure via Monte Carlo simulations in Section II.6. Then, we derive a mean-field equation for the dynamics of the average occupation, argue for why this description remains accurate, analyze the equation of state, and solve it in the homogeneous case in Section II.7. Lastly, we argue how these dual-unitary brick-wall circuits are advantageous compared to Clifford brick-wall circuits for the prediction of extensive observables using classical shadow tomography in Section II.8. This establishes our dual-unitary shadow tomography (DUST) protocol.
@@ II.1.1 General Formalism
is the locally-scrambled shadow norm $\|P\|_{\mathcal{E}_{\sigma}}^{2}$, which characterizes the sample complexity of predicting $P$ using the tomography scheme defined by $\mathcal{E}_{\sigma}$ [44, 42, 48].
@@ II.1.2 Dual-Unitarity in Qubits
$V(J)=\begin{pmatrix}e^{-iJ}&0&0&0\\ 0&0&-ie^{iJ}&0\\ 0&-ie^{iJ}&0&0\\ 0&0&0&e^{-iJ}\end{pmatrix}.$ (7)
At $J=\pi/4$, $V$ is proportional to a swap gate and simply interchanges sites, scrambling no information. As we turn $J$ away from $\pi/4$, we scramble more and more information until we reach $J=0$ which is akin to Hamiltonian generated scrambling with quantum $XY$-model. Additionally, the $J=0$ point is special since it is simply an iSWAP gate, which is known to be the fastest local scrambler [61]. As we will notice, this gate also generates the optimal dual-unitary shadow tomography scheme.
@@
In practice, it suffices to sample the local scramblers $u_{\pm},v_{\pm}$ according to the single-qubit Clifford group ${\mathcal{C}\ell}(2)$. This is because ${\mathcal{C}\ell}(2)$ forms a unitary 3-design over qubits, and this ensures Pauli-invariance, which in turn implies that the measurement channel is diagonal in the basis of Pauli strings [44].
II.2 Storing snapshot states
It is worth emphasizing that a key aspect of shadow tomography is the ability to store the snapshot states efficiently on a classical computer, for classical postprocessing and reconstructing the actual quantum state later. For the case of both global and shallow Clifford shadows, this follows from the Gottesman-Knill theorem, which allows us to store states of the form $U|b\rangle$ where $U$ is an $n$-qubit Clifford unitary and $|b\rangle$ is a computational basis state (corresponding to the bit string $b$). However, this method does not generalize to other classes of quantum circuits. Instead, for the case of dual-unitary shadows, this follows from the fact that we only ever need to use log-depth quantum circuits for shadow tomography. Namely, for predicting an operator with full-support, i.e., $\mathrm{supp}(O)=n$, we need a circuit depth $\sim\log(n)$. Such log-depth quantum circuits are simulable using standard matrix product state (MPS) methods [62] since we only need to simulate their evolution on computational basis states, $U|b\rangle$. In fact, this behavior is expected more generally for any locally-scrambled random quantum circuit, including Clifford shadows [51]. We discuss the optimality of log-depth quantum circuits for shadow tomography in Section II.8.
II.2.1 Set up
@@
II.3 Pauli Weight Transfer Matrix
@@
II.4 Dual-unitary case
@@
II.5 Pauli Weight Dynamics
@@ II.6 Occupation distribution
To refine our analysis, recall that the Pauli weight defines a probability distribution over all possible sub-regions of $\Omega=\{1\dots n\}$, or equivalently, all possible $n$-bit strings (see Eq. (18). This is because the transfer matrix which evolves the Pauli weight vector defined in Sec. II.2.1 is a probability-conserving process. Therefore, we have that the weights are positive i.e. $w_{P,t}(A):={\langle b|w_{P,t}\rangle}>0$, where $b$ is related to $A$ by Eq. (18), and also that
@@
II.7 Mean Field Approximation of Pauli Weight Dynamics
@@
II.8 Ramifications for the shadow norm and prediction
@@
which decays exponentially with a base of $e^{\log(25/16)}$, compared to $e^{\alpha^{\prime}}$ in the dual-unitary case. It follows that the dual-unitary should approach the Clifford limit more quickly provided that $\alpha^{\prime}>\log(25/16)$, which occurs around $\alpha^{*}_{MF}\approx 0.35$. However, we caution that in practice, it may require a very deep circuit to see the advantage for $\alpha<2/3$.
The “critical” value of $\alpha\approx 0.35$ makes sense when compared to studies of the rate of Local Operator Entanglement (LOE) growth in brick-wall dual-unitary circuits compared to brick-wall random (Haar) circuits [54, 70]. In these works, the authors investigate linear growth in the LOE as a signature of quantum chaos. In dual-unitary brick-wall circuits, the LOE velocity exceeds the velocity in the Haar (Clifford) case at $\alpha^{*}\approx 0.32$, which is remarkably close to the $\alpha^{*}_{MF}$. In appendix B, we show that the saturation rate for $\beta$ in the mean-field approximation is proportional to the local operator entanglement.
@@ IV Acknowledgements
Appendix A Mean-field derivation of the shadow norm bound